# Real-time 3D semantic occupancy prediction for autonomous vehicles using memory-efficient sparse convolution

### v1

@@ I INTRODUCTION
The rise of Autonomous Vehicles (AV) heralds a transformative era in transportation, moving us to potentially more efficient and safer alternatives. At the heart of these autonomous systems is a software stack composed of perception, localization, planning, and control. Among these, perception is crucial as it enables the vehicle to interpret and understand its surroundings, serving as the foundation for all subsequent vehicle decision-making. Typically, AV’s deploy a variety of sensors such as cameras and LiDARs in order to perceive its surroundings. Cameras operate in 2D perspective image space, capturing essential color and texture details but lacks 3D spatial information. This contrasts with sensors like LiDAR that inherently provide 3D coordinates, albeit often with lower spatial resolution. For a holistic 3D scene understanding, both pixel-level information from cameras and sparse spatial information from LiDARs should be leveraged.
@@
Traditional scene understanding uses 3D object detection methods that take in sensor information from camera and LiDARs and project 3D bounding boxes across objects of interest such as vehicles, pedestrians and traffic light. However, while 3D object detection provides information about object instances, it often fails to capture the complex object geometries and background information such as drivable space and buildings [1, 2]. An alternative approach to scene understanding is constructing 3D semantic occupancy maps. Given a physical distance from the ego vehicle in length, width and height, 3D semantic occupancy map partitions the environment into a structured grid map with a predefined resolution, allowing each grid cell to be assigned a semantic label.
Nevertheless, 3D semantic occupancy prediction of outdoor scenes proves to be difficult due to the scenes’ sparsity. Large areas such as the sky and occluded space captured by the camera often contain no relevant information, and most autonomous vehicle uses rotary beam LiDARs which are inherently sparse, leading to an overall sparse data. To accurately interpret and fill in the missing parts of the environment, these algorithms must incorporate elements of 3D scene completion in addition to 3D semantic segmentation.
Thereby, our main contributions are listed as the following:
@@
  • We evaluate the model’s 3D scene completion and semantic segmentation performance, achieving competitive accuracy against benchmark datasets [2, 1].
@@ II-A Camera View Transformation
Simple-BEV [16] suggests that the accuracy of Bird’s Eye View (BEV) and similarly, 3D prediction, often hinges more on factors like batch size, image resolution, and sensor fusion rather than sophisticated view transformation algorithms. Moreover, it is important to acknowledge that transforming views from 2D to 3D is inherently an ill-posed problem due to the absence of depth information in 2D images. This limitation highlights the necessity of employing deterministic methods that rely on additional sensors such as LiDARs, which are commonly included in most autonomous vehicles anyways. Arguing against the use of tools like LiDAR due to their cost overlooks their essential role in providing a fast and convenient method for accurate view transformation.
@@
More recently, multi-modal datasets such as nuScenes [23] and Occupancy 3D (Occ3D) [1] have enabled the development of vision-centric 3D semantic scene completion, primarily using cameras as the model input. Vision-centric methods, initially explored by Tesla [24] and further researched in various research studies [11, 1], employs cross-attention to geometrically relate 2D camera features into a 3D voxel grid, which is then used for BEV predictions or 3D semantic occupancy predictions. Efforts like deformable attention [14] and coarse-to-fine strategies [1] have been implemented to reduce the computational burden. Yet, a significant downside remains in the inherent dimensionality of establishing a dense 3D voxel grid, where attention mechanisms and transformers struggle to manage efficiently. Additionally, relying solely on camera data requires the model to simultaneously address the task of identifying and denoising distorted occluded areas on top of scene completion and semantic segmentation, posing a complex and challenging training task.
@@ III Problem Formulation
Formally, we define the 3D semantic occupancy prediction within the context of one front-view camera positioned to view directly ahead of an autonomous vehicle. Given a 2D camera image of dimensions $W\times H$, where each pixel $i_{ij}$ contains RGB values $(R,G,B)$ , the objective is to use view transformation to project and sparse convolution to densify and segment RGB and higher level image features into a 3D voxel space $V$ with dimensions $X\times Y\times Z$ in real-time. We define real-time to be at least 20 frames per second (20 FPS), with the goal of reaching 30 FPS to match human perception rate. Each voxel $v_{xyz}$ in this 3D space should encapsulate semantic and spatial information based on camera and LiDAR. The expected outcome is a labeled 3D voxel grid, where each voxel is affiliated with one of $L$ class labels, denoted as $L=\{l_{1},l_{2},...,l_{C}\}$, with $C$ representing the aggregate number of distinct classes.
@@ IV-A System Pipeline
The system pipeline is shown in Figure 2. Using a calibrated extrinsic matrix, $E_{\text{calib}}$ we first project LIDAR points onto the RGB image space. Each projected point is fused with RGB color information through bilinear interpolation. Subsequently, the fused data is unprojected back into the 3D space using the intrinsic camera parameters and the inverse of $E_{\text{calib}}$. We also perform feature extraction using EfficientNetV2 [27] at layer 3,4 and 6 to extract higher level 2D camera features and lift them to the 3D space using a the same projection method but scaled accordingly to accommodate for a reduction in spatial dimensions. The deterministic nature of the LiDAR-to-camera projection, assuming the sensors are adeptly calibrated, offers an immediate and precise depth-location correspondence - a key attribute for real-time tasks in autonomous vehicles.
Figure 2: Overview of System Pipeline
Initial representation now has attributes of spatial coordinates ($x,y,z$) and feature values ($[feats]+intensity$). Spatial coordinates is voxelized by a resolution of 0.4m, discretizing the pointcloud into a 3D tensor, $T_{\text{sparse}}$. $T_{\text{sparse}}$ has dimensions of 200 x 200 x 16 grid units in tensor shape, covering a physical boundary of [-40m, -40m, -1m] to [40m, 40m, 5.4m]. $T_{\text{sparse}}$ is later converted to Coordinate List format (COO) with only non-empty grids included during sparse convolution operations. We leverage the Minkowski Engine generative completion algorithm [28] to densify the sparse tensor: $T_{\text{dense}}=\mathcal{M}(T_{\text{sparse}})$, where $\mathcal{M}$ represents the generative scene completion operation. Scene completion is guided by Occ3D [1] dataset ground truth labels with the grid resolution to match $T_{\text{sparse}}$’s dimension.
We carry out semantic segmentation on $T_{\text{dense}}$, directly supervised by ground truth semantic labels. Both the semantic segmentation and scene completion step is trained as a multi-task problem, with the semantic segmentation having an additional encoder-decoder structure of smaller size, predicting outputs over 18 semantic categories. The final 3D semantic occupancy prediction has 200 x 200 x 16 voxel grid with 0.4m resolution.
@@ IV-C Network Architecture
With reference to figure 3, the proposed 3D semantic occupancy prediction model utilizes 2 U-Net-like [29] like encoder-decoder networks sequentially. Both U-Net shares similar structure, with the scene completion U-Net having deeper layers and specialized modules such as Squeeze and Excite [30], Generative Convolution Transpose and Pruning. An encoder block contains a double convolution Layer and a squeeze and excite (SE) layer. The convolution layer features batch normalization (BN) and ReLU activation to standardize the variable density of sparse data. The SE layer dynamically recalibrates channel-wise features through global average pooling, enhancing the model’s focus on key features across the input sparse data. For certain encoder blocks, image features extracted from the feature extractor are also concatenated before further downsampling.
The decoder module for scene completion is designed for generative tasks [28]. The process begins with generative transposed convolution capable of generating new coordinates with the outer product of the weight kernel and the input coordinates. A SE layer is attached after generation. At each decoder layer, a skip connection from the corresponding encoder is also added. Following that, a 3D binary classification convolution is applied to the sparse tensor which assigns a decision value to prune. The decision threshold is trainable with ground truth supervision. Redundant features are pruned using Minkowski Engine’s built in pruning module. Overall, the decoder generates, upsamples, classifies, and then prunes voxels to iteratively create accurate 3D representation at each layer depth.
Upon reaching the final decoder layer, an additional semantic segmentation U-Net head is attached for the sole purpose of semantic segmentation of the densified sparse tensor. This U-Net features convolution layers and skip connections, with bottleneck channel at 256.
@@ IV-D Loss Function
A multi-task loss function is defined to train the 3D Occupancy prediction model. The two terms are balanced by a Lambda ($\lambda$) constant empirically set to 0.5. Experiments conducted on different $\lambda$ constants is represented in the appendix.
@@
In normal semantic segmentation settings, multi-class cross entropy loss is sufficient with $z$ being predicted logits and $y$ being the ground truth labels. However, in nuScenes dataset, the class imbalance is extreme. For example, bicycle and motorcycle class is $10^{4}$ times lower in frequency than drivable space and free class [31]. Therefore, a class-balanced loss is applied [32], where an effective number $\beta$ is used to reweight the cross entropy loss function, giving more weight to classes with fewer effective samples. Effective number is set to $\beta=0.9$. Effective samples $\beta^{n_{y}}$ is calculated by the normalized number of voxels representing each class over all ground truth voxels over the entire training dataset. Validity of the class balanced loss is evaluated in the ablation study.
@@ V-A Experimental Setup
We utilize the Occupancy 3D (Occ3D) Dataset [1], an extension of the nuScenes dataset for 3D semantic occupancy benchmarking. It comprises of 28,130 Front camera RGB training images and 6019 validation images, as well as their corresponding voxelized ground truth semantic occupancy grid of size (200x200x16, 0.4m). The input LIDAR scans are obtained from the nuScenes [23] dataset. The training was configured with the Adam optimizer at a learning rate of 1e-4, and a Cosine-Annealing learning rate schedule. The model was trained till convergence, with a batch size of 10. The hardware setup for the training included 1 NVIDIA RTX 4090 GPU and 24GB DDR4 RAM. Training VRAM usage is at 4.3 GB.
Data augmentation is performed on the input and ground truth dataset, which includes RGB and LiDAR intensity value perturbation to simulate sensor noise (noise strength of 0.05 on normalized RGB values), random translation (-4 to 4 voxel grids), random yaw rotation (+/- 45 degrees), and ground truth voxel masking (5 percent).
V-B Scene Completion
We evaluate the 3D scene completion performance using binary voxel completion intersect over union (IoU), precision, recall, and F1 score to be compared with other scene completion methods. Referring to Table I, $P$ represents the set of voxels predicted by the model, $GT$ the set of voxels in the dense occupancy ground truth.
TABLE I: Evaluation metrics for 3D scene completion task.
Metric Equation Better
IoU $\frac{|P\cap GT|}{|P\cup GT|}$ Higher
Precision $\frac{|P\cap GT|}{|P|}$ Higher
Recall $\frac{|P\cap GT|}{|GT|}$ Higher
TABLE II: Quantitative results of 16 semantic classes on nuScenes validation dataset.
@@ V-B Scene Completion
Occ3D [1] (%) 28.53 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.0
Figure 4: Qualitative predictions on nuScenes validation dataset. Semantic occupancy grid is viewed in 3rd person perspective angle. Each group of images consists of occupancy prediction image on the left, ground truth label from Occ3D [1] on the right, and front camera image on top.
We compare these metrics with other scene completion methods on the nuScenes validation dataset. In particular, TransformerFusion [34] and Atlas [35] originally targets the task of 3D reconstruction of indoor scenes from RGB-D images. SurroundOcc [36], MonoScene [37] and Occ3D [1] focuses on the same outdoor driving scene completion task. Values from other methods are achieved through dense ground truth supervision for fair comparison, as reported by [36]. Referring to Table III, our proposed method achieves an scene completion IoU of 0.533, outperforming the current state-of-art by 16%. We also show a large improvement in precision, recall and f1, indicating we densify much more of the sparse input scene while being accurate to the ground truth.
TABLE III: Scene Completion evaluation on nuScenes validation set.
@@ V-C Semantic Segmentation
In semantic segmentation, mIoU (mean Intersection over Union) is calculated using the following formula: $\text{mIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$, where $N$ represents the total number of semantic classes, $TP_{i}$ denotes True Positives for class $i$, $FP_{i}$ represents False Positives for class $i$, and $FN_{i}$ corresponds to False Negatives for class $i$. We monitored the semantic segmentation performance through mean Intersection over Union (mIoU) and semantic IoU for each semantic class, and compare it with baseline BEVFormer [11], Occ3D [1], and NVOCC [33], current state-of-the-art. We do not compare the void and free class. Referring to Table II, we observe a 8 - 10 % improvement in mIoU over BEVFormer and Occ3D, as well as improvements in most semantic class IoU. Compared to NVOCC, we are competitive only in car, drivable surface, manmade and vegetation. However, bicycle, motorcycle and traffic cones presents a significant challenge to the model’s prediction. These small objects not only appear a lot less in Occ3D’s and nuScenes’s dataset, but also might be smoothed in the convolution processes, erasing or blending their features with its surroundings.
Figure 4 shows qualitative inferences of the model. Left group of day-time images demonstrates the model’s ability to classify scene objects of varying distance and shapes into distinct semantic classes. Right group of images shows model’s fidelity in low light environments: it is able to discern dark colored vehicles at night, as well as predicting obstacles on the left side of the drivable surface. However, the model does show a decrease in maximum prediction distance. For more qualitative results on different scenes, please refer to the appendix.

### v2

@@ I INTRODUCTION
The rise of Autonomous Vehicles (AV) heralds a transformative era in transportation, moving us to potentially more efficient and safer alternatives. At the heart of these autonomous systems is a software stack composed of perception, localization, planning, and control. Among these, perception is crucial as it enables the vehicle to interpret and understand its surroundings, serving as the foundation for all subsequent vehicle decision-making. Typically, AV’s deploy a suite of sensors such as cameras and LiDARs to perceive their surroundings. Cameras operate in 2D perspective image space, capturing essential color and texture details but lacking 3D spatial information. In contrast, sensors like LiDAR inherently provide 3D coordinates, albeit often with lower spatial resolution. For a holistic 3D scene understanding, both pixel-level information from cameras and sparse spatial information from LiDARs should be leveraged.
@@
Traditional scene understanding uses 3D object detection methods that take in sensor information from cameras and LiDARs and project 3D bounding boxes across objects of interest such as vehicles, pedestrians, and traffic lights. However, while 3D object detection provides information about object instances, it often fails to capture complex object geometries and background information such as drivable space and buildings [1, 2]. An alternative approach to scene understanding is constructing 3D semantic occupancy maps, which partition the environment into a structured grid map with a predefined resolution, allowing each grid cell to be assigned a semantic label.
Nevertheless, 3D semantic occupancy prediction of outdoor scenes proves to be difficult due to the scenes’ sparsity. Large areas such as the sky and occluded space captured by the camera often contain no relevant information. Most autonomous vehicles use rotary beam LiDARs which are inherently sparse, leading to overall sparse LiDAR data. To accurately interpret and fill in the missing parts of the environment, these algorithms must incorporate elements of 3D scene completion in addition to 3D semantic segmentation.
In this work, we focus on applying the existing sparse 3D convolution engine, Minkowski Engine [3], to a fused sensor representation of 2D camera and 3D LiDAR data for 3D semantic occupancy prediction. While the Minkowski Engine itself is a mature method, our key contribution lies in the innovative application of this sparse convolution framework to the problem of 3D semantic occupancy prediction. The fused sensor data enables sparse convolution to effectively handle outdoor scenes, allowing for accurate scene completion and semantic segmentation. Thereby, our main contributions are listed as the following:
@@
  • We evaluate the model’s 3D scene completion and semantic segmentation performance, achieving competitive accuracy against other algorithms on nuScenes dataset [1].
@@ II-A Camera View Transformation
Simple-BEV [16] suggests that the accuracy of Bird’s Eye View (BEV) semantic prediction and similarly, 3D semantic occupancy prediction, often hinges more on factors like batch size, image resolution, and sensor fusion rather than sophisticated view transformation algorithms. Moreover, it is important to acknowledge that transforming views from 2D to 3D is inherently an ill-posed problem due to the absence of depth information in 2D images. This limitation highlights the necessity of employing deterministic methods that rely on additional sensors such as LiDARs. Arguing against the use of tools like LiDAR due to their cost overlooks their essential role in providing a fast and accurate method for view transformation.
@@
More recently, multi-modal datasets such as nuScenes [23] and its extension, Occ3D-nuScenes [1], have enabled the development of vision-centric 3D semantic scene completion, primarily using cameras as the model input. Vision-centric methods, initially explored by Tesla [24] and further researched in various research studies [11, 25, 26, 27], employs cross-attention to geometrically relate 2D camera features into a 3D voxel grid, which is then used for BEV predictions or 3D semantic occupancy predictions. Efforts like deformable attention [14] and coarse-to-fine strategies [1] have been implemented to reduce the computational burden. Yet, a significant downside remains in the inherent dimensionality of establishing a dense 3D voxel grid, where attention mechanisms and transformers struggle to manage efficiently. Additionally, relying solely on camera data requires the model to simultaneously address the task of identifying and denoising distorted occluded areas on top of scene completion and semantic segmentation, posing a complex and challenging training task.
@@ III Problem Formulation
Formally, we define the 3D semantic occupancy prediction within the context of one front-view camera positioned to view directly ahead of an autonomous vehicle. Given a 2D camera image of dimensions $W\times H$, where each pixel $i_{ij}$ contains RGB values $(R,G,B)$ , the objective is to use view transformation to project, and sparse convolution to densify and segment image features into a 3D voxel space $V$ with dimensions $X\times Y\times Z$ in real-time. We define real-time to be at least 20 frames per second (20 FPS), with the goal of reaching 30 FPS to match human perception rate. Each voxel $v_{xyz}$ in this 3D space should encapsulate semantic and spatial information based on camera and LiDAR. The expected outcome is a labeled 3D voxel grid, where each voxel is affiliated with one of $L$ class labels, denoted as $L=\{l_{1},l_{2},...,l_{C}\}$, with $C$ representing the aggregate number of distinct classes.
@@ IV Method
We utilize the Occ3D-nuScenes dataset [1], an extension of the nuScenes dataset, which is designed for 3D semantic occupancy prediction and benchmarking. It comprises of 28,130 Front camera RGB training images and 6019 validation images, as well as their corresponding voxelized ground truth semantic occupancy grid. The input LIDAR scans are obtained from the original nuScenes [23] dataset.
@@ IV-A System Pipeline
The system pipeline is shown in Figure 2. Using a calibrated extrinsic and intrinsic matrix, we first project LIDAR points onto the RGB image space. Each projected point is fused with RGB color information through bilinear interpolation. Subsequently, the fused data is unprojected back into the 3D space. We also perform feature extraction using EfficientNetV2 [30] at layer 3,4 and 6 to extract higher level 2D camera features and lift them to the 3D space using the same projection method but scaled accordingly to accommodate for a reduction in spatial dimensions. The deterministic nature of the LiDAR-to-camera projection, assuming the sensors are adeptly calibrated, offers an immediate and precise depth-location correspondence - a key attribute for real-time tasks in autonomous vehicles.
Initial representation now has attributes of spatial coordinates ($x,y,z$) and feature values ($[feats]+LiDARintensity$). Spatial coordinates is voxelized by a resolution of 0.4m, discretizing the pointcloud into a 3D tensor, $T_{\text{sparse}}$. $T_{\text{sparse}}$ has dimensions of 200 x 200 x 16 grid units in tensor shape, covering a physical boundary of [-40m, -40m, -1m] to [40m, 40m, 5.4m]. $T_{\text{sparse}}$ is later converted to Coordinate List format (COO) with only non-empty grids included during sparse convolution operations. We leverage the Minkowski Engine generative completion algorithm [31] to densify the sparse tensor: $T_{\text{dense}}=\mathcal{M}(T_{\text{sparse}})$, where $\mathcal{M}$ represents the generative scene completion operation. Scene completion is guided by Occ3D-nuScenes [1] dataset ground truth labels with the grid resolution to match $T_{\text{sparse}}$’s dimension.
We carry out semantic segmentation on $T_{\text{dense}}$, directly supervised by ground truth semantic labels. Both the semantic segmentation and scene completion step is trained as a multi-task problem, with the semantic segmentation having an additional network of smaller size, predicting outputs over 18 semantic categories. The final 3D semantic occupancy prediction has 200 x 200 x 16 voxel grid with 0.4m resolution.
@@ IV-C Network Architecture
With reference to figure 3, the proposed 3D semantic occupancy prediction model utilizes 2 U-Net-like [32] like encoder-decoder networks sequentially. Both U-Net shares similar structure, with the scene completion U-Net having deeper layers and specialized modules such as Squeeze and Excite [33], Generative Convolution Transpose [31] and Sparse Pruning. An encoder block contains a double convolution Layer and a squeeze and excite (SE) layer. The convolution layer features batch normalization (BN) and ReLU activation to standardize the variable density of sparse data. The SE layer dynamically recalibrates channel-wise features through global average pooling, enhancing the model’s focus on key features across the input sparse data. For certain encoder blocks, image features extracted from EfficientNetV2 [30] are also concatenated before further downsampling.
The decoder module for scene completion is designed for generative tasks. The process begins with generative transposed convolution capable of generating new coordinates with the outer product of the weight kernel and the input coordinates. A SE layer is attached after generation. At each decoder layer, a skip connection from the corresponding encoder is also added. Following that, a 3D binary classification convolution is applied to the sparse tensor which assigns a decision value to prune. The decision threshold is trainable with ground truth supervision. Redundant features are pruned using Minkowski Engine’s built in pruning module. Overall, the decoder generates, upsamples, classifies, and then prunes voxels to iteratively create accurate 3D representation at each layer depth.
Upon reaching the final decoder layer, an additional semantic segmentation U-Net head is attached for the sole purpose of semantic segmentation of the densified sparse tensor. This U-Net features sparse convolution layers and skip connections, with bottleneck channel length of 256.
@@ IV-D Loss Function
A multi-task loss function is defined to train the 3D Semantic Occupancy prediction model. The two terms are balanced by a Lambda ($\lambda$) constant empirically set to 0.5. Experiments conducted on different $\lambda$ constants is represented in the ablation study.
@@
In normal semantic segmentation settings, multi-class cross entropy loss is sufficient with $z$ being predicted logits and $y$ being the ground truth labels. However, in Occ3D-nuScenes dataset, the class imbalance is extreme. For example, bicycle and motorcycle class is $10^{4}$ times lower in frequency than drivable space and free class [34]. Therefore, a class-balanced loss is applied [35], where an effective number $\beta$ is used to reweight the cross entropy loss function, giving more weight to classes with fewer effective samples. Effective number is set to $\beta=0.9$. Effective samples $\beta^{n_{y}}$ is calculated by the normalized number of voxels representing each class over all ground truth voxels over the entire training dataset. Validity of the class balanced loss is evaluated in the ablation study.
@@ V-A Experimental Setup
The training was configured with the Adam optimizer at a learning rate of 1e-4, and a Cosine-Annealing learning rate schedule. The model was trained till convergence, with a batch size of 10. The hardware setup for the training included 1 NVIDIA RTX 4090 GPU and 24GB DDR4 RAM. Training VRAM usage is at 4.3 GB.
Data augmentation is performed on the input and ground truth dataset, which includes image features and LiDAR intensity value perturbation to simulate sensor noise (noise strength of 5% on normalized values), random translation (-4 to 4 voxel grids) and ground truth voxel masking (10%).
TABLE I: Quantitative results of 16 semantic classes on Occ3D-nuScenes validation dataset.
@@
Figure 4: Qualitative predictions on Occ3D-nuScenes validation dataset. Semantic occupancy grid is viewed in 3rd person perspective angle. Each group of images consists of occupancy prediction image on the left, ground truth label on the right, and front camera image on top.
V-B Scene Completion
We evaluate the 3D scene completion performance using binary voxel completion intersect over union (IoU), precision, recall, and F1 score to be compared with other scene completion methods. Referring to Table II, $P$ represents the set of voxels predicted by the model, $GT$ the set of voxels in the dense occupancy ground truth.
TABLE II: Evaluation metrics for 3D scene completion task.
Metric Equation Better
IoU $\frac{|P\cap GT|}{|P\cup GT|}$ Higher
Precision $\frac{|P\cap GT|}{|P|}$ Higher
Recall $\frac{|P\cap GT|}{|GT|}$ Higher
TABLE III: Scene Completion evaluation on Occ3D-nuScenes validation set.
@@ V-B Scene Completion
We compare these metrics with other scene completion methods on the Occ3D-nuScenes validation dataset. In particular, TransformerFusion [38] and Atlas [39], which originally targets the task of 3D reconstruction of indoor scenes from RGB-D images. SurroundOcc [26], MonoScene [25] and CTF-Occ [1] focuses on the same outdoor driving scene completion task. Values from these methods are achieved through dense ground truth supervision for fair comparison, as reported by [26]. Referring to Table III, our proposed method achieves an scene completion IoU of 0.533, outperforming the current state-of-art by 16%. We also show a large improvement in precision, recall and f1, indicating we densify much more of the sparse input scene while being accurate to the ground truth.
@@ V-C Semantic Segmentation
In semantic segmentation, mIoU (mean Intersection over Union) is calculated using the following formula: $\text{mIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$, where $N$ represents the total number of semantic classes, $TP_{i}$ denotes True Positives for class $i$, $FP_{i}$ represents False Positives for class $i$, and $FN_{i}$ corresponds to False Negatives for class $i$. We monitored the semantic segmentation performance through mean Intersection over Union (mIoU) and semantic IoU for each semantic class, and compare it with baseline BEVFormer [11], several popular architectures [27, 1, 36] and NVOCC [37], current state-of-the-art. We do not compare the void and free class. Referring to Table I, we observe a 8 - 10 % improvement in mIoU over BEVFormer and CTF-Occ, as well as improvements in most semantic class IoU. Compared to NVOCC, we are competitive in car, drivable surface, manmade and vegetation. However, bicycle, motorcycle and traffic cones presents a significant challenge to the model’s prediction. These small objects not only appear a lot less in Occ3D-nuScene’s dataset due to class imbalance, but also potentially smoothed in the voxelization and convolution processes, erasing or blending their features with its surroundings. Methods to mitigate this problem will be discussed in section VI.
Figure 4 shows qualitative inferences of the model. Left group of day-time images demonstrates the model’s ability to classify scene objects of varying distance and shapes into distinct semantic classes. Right group of images shows model’s fidelity in low light environments: it is able to discern dark colored vehicles at night, as well as predicting obstacles on the left side of the drivable surface. However, the model does show a decrease in maximum prediction distance. More qualitative results are shown in Figure 5, 6, 7.